According to last year’s report, a filipino household spends an average of P400/day for food. Suppose you took a sample of 25 households. You determined how much each household spent for the food daily and the results revealed a mean of P380 and a standard deviation of P21.50. with 99% confidence what would be your conclusion



1
Expert's answer
2021-06-07T09:38:50-0400

"t=\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}=\frac{380-400}{21.5/\sqrt{25}}=-4.651"

"df=n-1=24"

Critical value:

"t_{.995}=2.797"

Since |t| > critical value, we can conclude that the claimed average of P400/day for food is not true.